A string of length L and mass M is lying on a horizontal table.A force F is applied at one of its ends.find the tension in the string at a distance x from the end at which the force is applied

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

 

Dear Rahul, 

acceleration of string =F/M, mass of portion of string between distance x and other end of string=M(1-x/L). tension at distance x=M(1-x/L)F/M=F(1-x/L)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras